MBI and Pear are the only two producers of computers. MBI started producing earlier than Pear. MBI's costs of production are given by C1(y1)=y12. Pear's cost function is C2(y2)=5y2. The national demand for computers is y=106-105p.

a) Calculate the Stackelberg equilibrium in which MBI is the leader in this market. Indicate output levels, market price, and the profits of each firm.

b) Suppose that both firms enter this market at the same time. Calculate the Cournot equilibrium and compare it to the situation in part (a).
Answer #1

Y2 *= 249998.125

Total Y = 250001.875

Now P* = 10 -( 250001.875)/105

P= 7.5

Now profit of first firm = 7.5*3.75 - (3.75)2

π1= 14.0625

π2 = (7.5-5)*249998.125

= 624,995.313 answer